# NavMoE short=NavMoE, long=Navigation via Mixture of Experts

### v1

@@ I Introduction
Combining experts with different expertise is a effective approach in ML. Classical ensemble learning methods like [14] or Mixture of Experts (MoE) [15, 16] often leads to performance improvements. Recent works aim to address the above challenges by integrating model-based and learning-based methods into a unified framework. [17] encodes images and point clouds into a shared 3D map for geometric analysis and traversability prediction but depends heavily on high-quality multi-sensory data for training. [18] serially connect both paradigms using a human-designed probing strategy, but lacks cross-domain flexibility and is inefficient as both methods must always run. [19] utilizes physics-informed learning to improve generalization through simulation, but the domain gap remains a challenge. However, these hybrid approaches still face significant challenges for traversability estimation: 1) It is difficult to determine the contribution of each method, as their performance can vary widely across domains. 2) Hybrid networks often require extensive expensive synchronized multi-sensory data for training. 3) It is often difficult to jointly update the components to enable better cooperation among them. This is because learning-based methods rely on gradient-based optimization, while model-based methods are usually nondifferentiable.
@@ II Hierarchical and Modular Mixture-of-Experts
An overview of NavMoE is shown in Fig. 2. Given a RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ and a depth image $\mathbf{D}\in\mathbb{R}^{H\times W}$, NavMoE predicts a traversability map $\mathcal{T}\in\mathbb{R}^{H\times W}$ and then computes a feasible robot trajectory $\tau\in\mathbb{R}^{3}$. NavMoE introduces two levels of specialization: domain experts $\mathbf{E}_{n}^{\rm Domain}$ with $n=\{1,2,...,N\}$ at the top level and terrain experts $\mathbf{E}_{k}^{\rm Terrain}$ with $k=\{1,2,...,K\}$ within each domain. Notably, K is the number of terrain experts within each domain, so in our setup K=2 and N=3 as illustrated in Fig. 2 Domain experts route input images to relevant high-level domains for targeted optimization, while terrain experts further refine predictions based on distinct terrain features.
@@
where we use DINOv2 [20] as the encoder $\mathcal{E}$. After that, the top-level router $r^{\rm top}$ takes $\mathbf{h}$ for domain classification. The output of the top-level router is a normalized vector $\hat{\mathbf{w}_{n}}\in\mathbb{R}^{N\times 1}$, which indicates the predicted possibilities for each type of domain: $\hat{\mathbf{w}_{n}}=r^{\rm top}(\mathbf{h})$. Then, based on the predicted weights $\hat{\mathbf{w}_{n}}$, the domain experts $\mathbf{E}^{\rm Domain}$ are queried using the proposed lazing gating algorithm (detailed in Section III).
@@
Within the selected domain, we implement a lower-level expert routing mechanism. Specifically, a router $r^{\rm lower}$ assigns terrain experts $\mathbf{E}^{\rm Terrain}_{nk}$ to each pixel based on a pixel-wise terrain partition of the input image. Unlike the top-level coarse classification that assigns the whole input to several domain categories, $r^{\rm lower}$ performs fine-grained segmentation by predicting pixel-wise terrain expert preferences.It takes feature map $\mathbf{H}$ as inputs and predicts a weight map of expert weights $\hat{\mathbf{W}}\in\mathbb{R}^{H\times W\times K}$:
$\hat{\mathbf{W}}=r^{\rm lower}(\mathbf{H})$ (2)
@@
Each terrain expert $\mathbf{E}^{\rm Terrain}_{nk}$ is implemented as either a neural network or a model-based algorithm (details in Section V). These experts take RGB or depth images as input and produce pixel-aligned traversability maps $\mathcal{T}_{nk}$. Formally, an expert can be expressed as a mapping $E^{\rm Terrain}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H\times W}$, where $C=3$ if the input is RGB image, $C=1$ if the input is depth image. The final $\mathcal{T}$ is computed as a two-level weighted combination of expert predictions:
$\mathcal{T}=\sum_{u,v}^{I}\sum_{n}^{N}\left(\hat{\mathbf{w}}_{n}\cdot\sum_{k}^{K}\left(\hat{\mathbf{W}}_{k}^{\rm lower}[u,v]\cdot\mathcal{T}_{nk}[u,v]\right)\right).$ (3)
@@
The resulting $\mathcal{T}$ is then passed to a motion planner to compute a trajectory $\tau^{*}$. Here we adopt the primitive-based planner [6]. Define the candidate motion primitives set on $\mathcal{T}$ as $S(\mathcal{T})=\{\tau_{1},\tau_{2},...,\tau_{k}\}$, we formulate the trajectory optimization problem as:
$\tau^{*}=\operatorname*{arg\,min}_{\tau\in S(\mathcal{T})}\sum_{i=1}^{k}J_{trav}\left(\tau_{i},\mathcal{T}\right)+\lambda J_{dis}\left(\tau_{i},p_{goal}\right).$ (4)
@@
$J_{trav}\left(\tau_{i},\mathcal{T}\right)$ represents the traversability cost of the path $\tau_{i}$ on $\mathcal{T}$, $J_{dis}\left(\tau_{i},p_{goal}\right)$ represents the cost of $\tau_{i}$ to the goal $p_{goal}$.
@@
Specifically, for $J_{trav}\left(\tau,\mathcal{T}\right)$, let $J_{\text{trav}}(\tau,\mathcal{T})$ denote the traversability cost of a 3D trajectory $\tau$ with respect to a 2D traversability map $\mathcal{T}$. We first orthogonally project $\tau=\{q_{1},q_{2},\dots,q_{n}\}$, where $q_{i}\in\mathbb{R}^{3}$, onto the 2D plane of $\mathcal{T}$, yielding $\hat{\tau}=\{p_{1},p_{2},\dots,p_{n}\}$, where each $p_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$ corresponds to a waypoint on the map. The traversability cost is then computed by summing the map values at each projected waypoint:
$J_{\text{trav}}(\tau,\mathcal{T})=\frac{\sum_{p_{i}\in\hat{\tau}}\left(1-\mathcal{T}\left(p_{i}\right)\right)}{|\tau|}.$ (5)
@@
For $J_{dis}\left(\tau,p_{goal}\right)$, let $\tau=\{q_{1},q_{2},\dots,q_{n}\}$ denote a candidate 3D trajectory and $p_{\text{goal}}\in\mathbb{R}^{3}$ be the goal position. Let $p\in\mathbb{R}^{3}$ be the current robot position. Define the distance-based cost as:
$J_{\text{dis}}(\tau,p_{\text{goal}})=1-\frac{\|p-p_{\text{goal}}\|-\min_{q_{j}\in\tau}\|q_{j}-p_{\text{goal}}\|}{H},$ (6)
where $H$ denotes the planning horizon. $H$ is set as $2m$ for indoor, $4m$ for unstructured outdoor and $8m$ for unstructured outdoor environment. The result lies in the range of $[0,1]$.
@@ III-A Lazy Gating Mechanism
As the number of experts increase, activating all experts can cause severe computational overhead. To address this challenge, we introduce the lazy gating mechanism as detailed in Algorithm 1. The key idea is to selectively activating a subset of experts that could significantly impact path solution.
@@
2:  Compute $\hat{\mathbf{W}}\in\mathbb{R}^{H\times W\times K}$ using Eq. 2
@@
9:     $\mathcal{T}\leftarrow\hat{\mathcal{T}}$, $\mathbf{W}^{\rm prior}\leftarrow\hat{\mathbf{W}}[k]$
@@
11:     $\mathcal{T}\leftarrow\frac{\mathcal{T}\cdot\mathbf{W}^{\rm prior}+\hat{\mathcal{T}}\cdot\hat{\mathbf{W}}[k]}{\mathbf{W}^{\rm prior}+\hat{\mathbf{W}}[k]}\triangleright$ Eq. 8
12:     $\mathbf{W}^{\rm prior}\leftarrow\mathbf{W}^{\rm prior}+\hat{\mathbf{W}}[k]$
@@
notably, here $\mathcal{K}$ equals to $K$ defined above, we adopt a different notation to represent the reordering process used to construct the queue. Intuitively, if experts have similar weights, we prioritize those with higher computational efficiency; if one has a dominant weight, we prioritize even if it has lower efficiency. As shown in Fig. 3 A, the lazy gating algorithm processes experts sequentially, iterating from the front of the queue to the rear. During each iteration, the next expert is dequeued and queried to provide a traversability prediction. The internal processing of each expert is shown in Fig. 3 B: At iteration $k$, the algorithm maintains a prior traversability map $\mathcal{T}_{k-1}$, which aggregates the predictions from the previous k1k-1 experts. The $k$th expert predicts a traversability map $\hat{\mathcal{T}}_{k}$ while the gating network estimates the current expert’s confidence as weight $\mathbf{W}_{k}$. The current expert’s output $\hat{\mathcal{T}}_{k}$ is then fused into the overall map:
$\mathcal{T}_{k}=\left.\left(\mathcal{T}_{k-1}\cdot\sum_{j=1}^{k-1}\hat{\mathbf{W}}_{j}+\hat{\mathcal{T}}_{k}\cdot\hat{\mathbf{W}}_{k}\right)\right/\sum_{j=1}^{k}\hat{\mathbf{W}}_{j},$ (8)
@@
where the updated traversability map $\mathcal{T}_{k}$ is then used as the new prior for the next iteration. The planner will take $\mathcal{T}_{k}$ as input to compute the path $\tau_{k}$ and path cost $C(\tau_{k})$ using Eq. 4. Given $C(\tau_{k})$ as prior, the expected cost of the final path $C(\tau_{K})$ will be constrained to lie within a value interval after any admissible update from expert $k+1$ to expert $K$; we discuss how to get the value bounds in the following proposition. To reduce the computation overhead, we prune the rest of the experts if they do not significantly change the traversability output.
@@
Proposition (Convergence Bound on Path Cost Under Incremental Traversability Updates). Let $\mathcal{T}_{k}$ be the traversability map after $k$ aggregative updates following Eq. 8, and let $C(\mathcal{T}_{k})\doteq C(\tau^{*})$ denote the path cost computed using $\mathcal{T}_{k}$. Define upper and lower expectation maps based on the remaining uncertainty in traversability ($[0,1]$) from step $k+1$ to $K$ as:
@@
$\underline{\mathcal{T}_{k}}$ $=\left.\left(\mathcal{T}_{k}\cdot\sum_{j=1}^{k}\hat{\mathbf{W}}_{j}+0\cdot\sum_{j=k+1}^{K}\hat{\mathbf{W}}_{j}\right)\right/\sum_{j=1}^{K}\hat{\mathbf{W}}_{j}$ (9a)
$=\mathcal{T}_{k}\cdot\sum_{j=1}^{k}\hat{\mathbf{W}}_{j}\left/\sum_{j=1}^{K}\hat{\mathbf{W}}_{j}\right$.,
@@
$\overline{\mathcal{T}_{k}}=\left.\left(\mathcal{T}_{k}\cdot\sum_{j=1}^{k}\hat{\mathbf{W}}_{j}+1\cdot\sum_{j=k+1}^{K}\hat{\mathbf{W}}_{j}\right)\right/\sum_{j=1}^{K}\hat{\mathbf{W}}_{j},$ (9b)
@@
where $\hat{\mathbf{W}}_{j}$ denotes the weight related to the $j$-th update predicted by the routers, and $K$ is the total number of updates. Intuitively, the lower expectation map assumes that all remaining (unexecuted) experts label the environment as untraversable, while the upper expectation map assumes they label it as fully traversable. Then, the deviation of the current path cost from the final one is bounded as

### v2

@@ I Introduction
Combining experts with different expertise is a effective approach in ML. Classical ensemble learning methods like [14] or Mixture of Experts (MoE) [15, 16] often leads to performance improvements. Recent works aim to address the above challenges by integrating model-based and learning-based methods into a unified framework. [17] encodes images and point clouds into a shared 3D map for geometric analysis and traversability prediction but depends heavily on high-quality multi-sensory data for training. [18] serially connect both paradigms using a human-designed probing strategy, but lacks cross-domain flexibility and is inefficient as both methods must always run. [19] utilizes physics-informed learning to improve generalization through simulation, but the domain gap remains a challenge. However, these hybrid approaches still face significant challenges for traversability estimation: 1) It is difficult to determine the contribution of each method, as their performance can vary widely across domains. 2) Hybrid networks often require extensive expensive synchronized multi-sensory data for training. 3) It is often difficult to jointly update the components to enable better cooperation among them. This is because learning-based methods rely on gradient-based optimization, while model-based methods are usually nondifferentiable, as they often involve discrete decision rules or search-based operations that break gradient flow.
@@ II Hierarchical and Modular Mixture-of-Experts
An overview of NavMoE is shown in Fig. 2. Given a RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ and a depth image $\mathbf{D}\in\mathbb{R}^{H\times W}$, NavMoE predicts a traversability map $\mathcal{T}\in\mathbb{R}^{H\times W}$ and then computes a feasible robot trajectory $\tau\in\mathbb{R}^{3\times l}$, where $l$ is the trajectory length. NavMoE introduces two levels of specialization: domain experts $\mathbf{E}_{m}^{\rm Domain}$ with $m=\{1,2,...,M\}$ at the top level and terrain experts $\mathbf{E}_{mn}^{\rm Terrain}$ with $n=\{1,2,...,N\}$ within each domain $m$. Notably, $N$ is the number of terrain experts within each domain; in our setup, $M=3$ and $N=2$, as illustrated in Fig. 2. Domain experts route input images to relevant high-level domains for targeted optimization, while terrain experts further refine predictions based on distinct terrain features.
@@
where we use DINOv2 [20] as the encoder $\mathcal{E}$. After that, the top-level router $r^{\rm top}:\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}^{M}$ takes $\mathbf{h}$ as input for domain classification. The output of the top-level router is a normalized vector $\mathbf{w}_{m}\in\mathbb{R}^{M\times 1}$, which indicates the predicted possibilities for each type of domain: $\mathbf{w}_{m}=r^{\rm top}(\mathbf{h})$. Then, based on the predicted weights $\mathbf{w}_{m}$, the domain experts $\mathbf{E}^{\rm Domain}$ are queried using the proposed lazy gating algorithm (detailed in Section III).
@@
Within the selected domain, we implement a lower-level expert routing mechanism. Specifically, a router $r^{\rm lower}$ assigns terrain experts $\mathbf{E}^{\rm Terrain}_{mn}$ to each pixel based on a pixel-wise terrain partition of the input image. Unlike the top-level coarse classification that assigns the whole input to several domain categories, $r^{\rm lower}:\mathbb{R}^{H\times W\times d_{1}}\rightarrow\mathbb{R}^{H\times W\times N}$ performs fine-grained segmentation by predicting pixel-wise terrain expert preferences. It takes feature map $\mathbf{H}$ as input and predicts a weight map of expert weights $\mathbf{W}\in\mathbb{R}^{H\times W\times N}$:
$\mathbf{W}=r^{\rm lower}(\mathbf{H})$ (2)
@@
Therefore, weight map for each expert can be expressed as $\mathbf{W}_{n}\in\mathbb{R}^{H\times W}$. Each terrain expert $\mathbf{E}^{\rm Terrain}_{mn}$ is implemented as either a neural network or a model-based algorithm (details in Section V). These experts take RGB or depth images as input and produce pixel-aligned traversability maps $\mathcal{T}_{mn}\in\mathbb{R}^{H\times W}$. Formally, an expert can be expressed as a mapping $\mathbf{E}_{mn}^{\rm Terrain}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H\times W}$, where $C=3$ if the input is an RGB image and $C=1$ if the input is a depth image. The final $\mathcal{T}$ is computed as a two-level weighted combination of expert predictions:
$\mathcal{T}=\sum_{m=1}^{M}\left(\mathbf{w}_{m}\cdot\sum_{n=1}^{N}\left(\mathbf{W}_{n}^{\rm lower}\odot\mathcal{T}_{mn}\right)\right),$ (3)
@@
where $\odot$ refers to elementwise multiplication. The resulting $\mathcal{T}$ is then passed to a motion planner to compute an optimal trajectory $\tau^{*}$. Here we adopt the primitive-based planner in [6]. Defining the candidate motion primitives set for the traversability map $\mathcal{T}$ as $S(\mathcal{T})=\{\tau_{1},\tau_{2},...,\tau_{l}\}$, we formulate the trajectory optimization problem as:
$f(\mathcal{T},p_{\mathrm{goal}})\triangleq\operatorname*{arg\,min}_{\tau\in S(\mathcal{T})}\sum_{i=1}^{\ell}\Big{(}J_{\mathrm{trav}}(\tau_{i},\mathcal{T})+\lambda\,J_{\mathrm{dis}}(\tau_{i},p_{\mathrm{goal}})\Big{)}.$ (4)
@@
$J_{\textrm{trav}}\left(\tau_{i},\mathcal{T}\right)$ represents the traversability cost of the path $\tau_{i}$ on $\mathcal{T}$, $J_{\textrm{dis}}\left(\tau_{i},p_{\textrm{goal}}\right)$ represents the cost of $\tau_{i}$ with respect to reaching the goal $p_{\textrm{goal}}$. Finally, $\tau^{*}=f(\mathcal{T},p_{\mathrm{goal}})$.
@@
Specifically, let $J_{\text{trav}}(\tau,\mathcal{T})$ denote the traversability cost of a 3D trajectory $\tau$ with respect to a 2D traversability map $\mathcal{T}$. Utilizing the depth information, we first project $\tau=\{q_{1},q_{2},\dots,q_{l}\}$, where $q_{i}\in\mathbb{R}^{3}$, onto the 2D plane of $\mathcal{T}$, yielding $\hat{\tau}=\{p_{1},p_{2},\dots,p_{l}\}$, where each $p_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$ corresponds to a waypoint on the map. The traversability cost is then computed by summing the map values at each projected waypoint: $J_{\text{trav}}(\tau,\mathcal{T})=\left.\sum_{p_{i}\in\hat{\tau}}\left(1-\mathcal{T}\left(p_{i}\right)\right)\right/|\tau|.$ Notably, since $J_{\text{trav}}(\tau,\mathcal{T})$ is defined as a cost, lower values indicate better traversability.
@@
For $J_{\textrm{dis}}\left(\tau,p_{\textrm{goal}}\right)$, let $\tau=\{q_{1},q_{2},\dots,q_{l}\}$ denote a candidate 3D trajectory and $p_{\text{goal}}\in\mathbb{R}^{3}$ be the goal position. Let $p\in\mathbb{R}^{3}$ be the current robot position. Define the distance-based cost as: $J_{\text{dis}}(\tau,p_{\text{goal}})=\left.1-(\|p-p_{\text{goal}}\|-\min_{q_{j}\in\tau}\|q_{j}-p_{\text{goal}}\|)\right/H,$ where $H$ denotes the maximum planning distance, which means that for all $q_{j}$ in $\tau$, $\|q_{j}-p\|\leq H$. $H$ is set as $2m$ for indoor, $4m$ for unstructured outdoor and $8m$ for unstructured outdoor environment. The result lies in the range of $[0,1]$.
@@ III-A Lazy Gating Mechanism
As the number of experts increases, activating all experts can cause severe computational overhead. To address this challenge, we introduce the lazy gating mechanism as detailed in Algorithm 1. The key idea is to selectively activate a subset of experts that could significantly impact the path solution. We implement the lazy gating mechanism to both lower and higher level routers.
@@
2:  Compute $\mathbf{W}\in\mathbb{R}^{H\times W\times N}$ using Eq. 2
@@
9:     $\mathcal{T}\leftarrow\hat{\mathcal{T}}$, $\mathbf{W}^{\rm prior}\leftarrow\mathbf{W}[n]$
@@
11:     $\mathcal{T}\leftarrow\frac{\mathcal{T}\cdot\mathbf{W}^{\rm prior}+\hat{\mathcal{T}}\cdot\mathbf{W}[k]}{\mathbf{W}^{\rm prior}+\mathbf{W}[k]}\triangleright$ Eq. 6
12:     $\mathbf{W}^{\rm prior}\leftarrow\mathbf{W}^{\rm prior}+\mathbf{W}[k]$
@@
Here, $\mathcal{K}=K$ defined in Fig. 1; however, we adopt a different notation to represent the reordering process used to construct the queue. Intuitively, if experts have similar weights, we prioritize those with higher computational efficiency; if one has a dominant weight, we prioritize even if it has lower efficiency. As shown in Fig. 3 A, the lazy gating algorithm processes experts sequentially, iterating from the front of the queue to the rear. During each iteration, the next expert is dequeued and queried to provide a traversability prediction. The internal processing of each expert is shown in Fig. 3 B: At iteration $k$, the algorithm maintains a prior traversability map $\mathcal{T}_{k-1}$, which aggregates the predictions from the previous k1k-1 experts. The $k$th expert predicts a traversability map $\hat{\mathcal{T}}_{k}$ while the gating network estimates the current expert’s confidence as weight $\mathbf{W}_{k}$. The current expert’s output $\hat{\mathcal{T}}_{k}$ is then fused into the overall map:
$\mathcal{T}_{k}=\left.\left(\mathcal{T}_{k-1}\odot\textstyle\sum_{j=1}^{k-1}\mathbf{W}_{j}+\mathcal{T}_{k}\odot\mathbf{W}_{k}\right)\right/\textstyle\sum_{j=1}^{k}\mathbf{W}_{j},$ (6)
@@
where the division is performed elementwise. The updated traversability map $\mathcal{T}_{k}$ is then used as the new prior for the next iteration. The planner will take $\mathcal{T}_{k}$ as input to compute the path $\tau_{k}$ and path cost $C(\tau_{k})$ using Eq. 4. Given $C(\tau_{k})$ as prior, the expected cost of the final path $C(\tau_{K})$ will be constrained to lie within an interval after any admissible update from expert $k+1$ to expert $K$; we discuss how to get the bounds of this interval in the following proposition. To reduce the computation overhead, we prune the rest of the experts if they do not significantly change the traversability output by the threshold $\epsilon$, as detailed in lines 18-25 of Alg. 1.
Proposition 1 (Convergence Bound on Path Cost Under Incremental Traversability Updates)
@@
Let $\mathcal{T}_{k}$ be the traversability map after $k$ aggregative updates following Eq. 6, and let $C(\mathcal{T}_{k})\doteq C(\tau^{*})$ denote the path cost computed using $\mathcal{T}_{k}$. Define upper and lower expectation maps based on the remaining uncertainty in traversability ($[0,1]$) from step $k+1$ to $K$ as:
@@
$\underline{\mathcal{T}_{k}}$ $=\left.\left(\mathcal{T}_{k}\odot\textstyle\sum_{j=1}^{k}\mathbf{W}_{j}+0\cdot\textstyle\sum_{j=k+1}^{K}\mathbf{W}_{j}\right)\right/\textstyle\sum_{j=1}^{K}\mathbf{W}_{j}$ (7a)
$=\mathcal{T}_{k}\odot\textstyle\sum_{j=1}^{k}\mathbf{W}_{j}\left/\textstyle\sum_{j=1}^{K}\mathbf{W}_{j}\right$.,
@@
$\overline{\mathcal{T}_{k}}=\left.\left(\mathcal{T}_{k}\odot\textstyle\sum_{j=1}^{k}\mathbf{W}_{j}+1\cdot\textstyle\sum_{j=k+1}^{K}\mathbf{W}_{j}\right)\right/\textstyle\sum_{j=1}^{K}\mathbf{W}_{j},$ (7b)
@@
where $\mathbf{W}_{j}$ denotes the weight related to the $j$-th update predicted by the routers, and $K$ is the total number of updates. Intuitively, the lower expectation map assumes that all remaining (unexecuted) experts label the environment as untraversable, while the upper expectation map assumes they label it as fully traversable. Then, the deviation of the current path cost from the final one is bounded as
@@
Figure 4: $\mathcal{T}$ is modeled as an undirected graph $G=(V,E)$. A path is modeled as a sequence of vertices $\tau$ connected via $G$.